# The Influence of Code Smells in Efferent Neighbors on Class Stability

### v1

@@ 3.2.4. RQ4
We speculate that this situation may further influence stability. This case arises when the focal class and an efferent neighbor both contain code smells (i.e., efferent code smell coupling), and the two smell instances are directly connected by a static dependency where the focal class is the source artifact. In such cases, when a smell in the efferent neighbor changes, the dependency makes it more likely that developers must also change the corresponding smell in the focal class. As a result, developers are more likely to need to modify smell instances in two classes at once. In contrast, in efferent code smell coupling without a direct static dependency between the smell instances, the probability of needing to change smells in both classes at once is lower. Additionally, since interacting code smells are associated with more static dependencies between them (Zhang et al., 2025), there are more routes for changes to ripple into the focal class, making it more likely to be changed more often and more extensively. Hence, our RQ4 and its hypotheses are:
@@ 3.3. Independent Variables (IVs)
The rationale for measuring the number and variety of code smells is that higher values are often associated with degraded maintainability, particularly reduced stability (Palomba et al., 2018b; Santana et al., 2024b). The rationale for measuring interaction intensity is that higher intensity reflects a greater number of dependencies connecting smelly artifacts, which can harm comprehensibility and compound a developer’s ability to effectively perform modifications (Yamashita and Moonen, 2013).
@@ 3.5. Control Variables (CVs)
The rationale for including these controls is to reduce confounding. Larger classes inherently have a higher probability of being changed (Olbrich et al., 2010), and classes with more efferent neighbors are also more likely to change because modifications to any of their efferent neighbors may trigger changes in the focal class (Martin, 1994, 2000).
@@ 3.7. System Selection
We will conduct our study on top-starred open-source projects from GitHub. The rationale for using top-starred projects is to prioritize the most popular and widely adopted projects (Santana et al., 2024b). By ranking projects in descending order of the number of stars, we will select the first 100 projects that meet the following inclusion criteria (IC), following recommendations from Kalliamvakou et al. (Kalliamvakou et al., 2016): IC1. More than 100 forks (indicates meaningful developer interest). IC2. At least 20 contributors (to avoid personal or low-activity projects). IC3. At least one year of commit history with at least 50 commits within that one-year period (to ensure meaningful development activity, approximately one commit per week). IC4. More than 80% of the code written in Java (due to the smell detection tool limitation). IC5. Not an educational project (educational projects may not represent real software development processes).
@@ 3.8. Data Analysis
It is reported that most software engineering metrics do not follow a normal distribution (Beranič and Heričko, 2020). Therefore, we assume our IVs, DVs, and CVs are non-normal. We use statistical methods that do not assume normality and apply log transformations when appropriate. Due to the large-scale nature of our dataset, normality tests become overly sensitive; therefore, even though we assume non-normality, we verify practical deviations by examining quantile-quantile plots.
@@
We analyze our DVs using generalized linear models (GLMs). GLMs extend ordinary linear regression to non-normal outcomes by using an appropriate distribution and a link function to relate predictors to the expected outcome (McCullagh, 2019). In empirical software engineering, regression-based models such as GLMs are widely used to model a variety of software engineering data (Canfora et al., 2015; Ahmed et al., 2017). In our analysis, for each hypothesis, we fit a separate GLM with the corresponding DV (either ChF or ChS) as the response and the IV and CVs specified for that hypothesis in Table 2 as predictors.
@@
Both DVs are count data (i.e., non-negative integers arising from counting) and are modeled using negative binomial GLMs with a log link. Poisson and negative binomial GLMs are both standard for count data, but Poisson assumes equidispersion (the mean equals the variance) (McCullagh, 2019). Because previous work suggests software engineering count measures can be over-dispersed relative to this assumption, we adopt the negative binomial model for its flexibility (Ostrand et al., 2004). For completeness, we also fit Poisson GLMs and compute the dispersion statistic to confirm the presence of overdispersion.
@@
Since we assume non-normality, we apply log transformations to our count-based IVs (ClSize, #EffNei, #SmellFoc, #SmellEff, VarSmellFoc, VarSmellEff) before modeling, to reduce the influence of extreme values and obtain smoother relationships (Changyong et al., 2014). We do not transform the DVs, as they are modeled directly using appropriate GLMs.
@@
A random effect is a model component that captures group-level variability, and it is increasingly used in empirical software engineering (Gelman and Hill, 2007; Kuutila et al., 2021; Lambiase et al., 2024). We include a project-level random intercept in all models to account for systematic differences between projects, such as variation in age, development activity, commit practices, and team size. This adjustment ensures that comparisons between classes are not biased by project-specific factors while allowing the estimated effects of code smells to generalize across projects (Raudenbush and Bryk, 2002). Using a random intercept rather than fixed effects is appropriate given the large number of projects and our aim to estimate population-level effects that generalize beyond the specific projects, rather than fitting a separate parameter for each project (Gelman and Hill, 2007).
@@
For each fitted model, we will report the regression coefficient ($\beta$) for each IV together with its standard error, 95% confidence interval, and $p$-value (computed using Wald tests for the regression coefficients; one-sided as specified by each directional hypothesis), which collectively describe the estimated direction of the effect, its precision, and its statistical significance. To convey the practical magnitude of effects (i.e., effect size), we will report Incidence Rate Ratios (IRRs) (Hilbe, 2011) and Average Marginal Effects (AMEs) (Long and Freese, 2006). These translate the estimated coefficients into interpretable changes in complementary ways: IRRs present effects as percentage (multiplicative) changes, while AMEs present effects as average raw-unit (additive) changes. Model performance will be summarized using the log-likelihood (McCullagh, 2019) and McFadden’s pseudo-$R^{2}$ (McFadden, 1972), both of which provide standard goodness-of-fit evaluations for GLMs. Together, these reporting elements ensure a comprehensive assessment of model adequacy.
@@ 3.9. Implementation
Projects will be filtered (according to the first four inclusion criteria discussed in Section 3.7), selected, and downloaded automatically using the GitHub API (17). Inclusion Criterion 5 (i.e., not an educational project) will be enforced manually by inspecting each project repository, primarily the project documentation (e.g., the README). Code smells will be detected using JSpIRIT (S. Vidal, H. Vazquez, J. A. Diaz-Pace, C. Marcos, A. Garcia, and W. Oizumi (2015); 23). The rationale for choosing this tool is that it can detect the ten code smells we will study, and it strictly follows the definitions and detection strategies provided by Lanza and Marinescu (Lanza and Marinescu, 2007), ensuring transparency. Static dependencies, efferent neighbors of the focal class, and class size will be detected and measured using the static analysis tool CodeQL (10). This information will be collectively analysed using Python to obtain all IVs and CVs. Git commands (16) will be used to analyse commit history and extract information such as diffs, and RefactoringMiner (45) will be used to identify class refactorings, including class rename, move, split, and merge. Python will automate the process and compute the DVs for each included class. Finally, R will be used to fit the GLMs via the glmmTMB package (18) and to obtain all statistical outputs.
@@ 4. Threats to Validity
A further threat to internal validity concerns our project selection process. Following the recommendations of Kalliamvakou et al. (Kalliamvakou et al., 2016), we exclude personal, inactive, and educational repositories and focus on widely adopted projects to maximize insights into real code evolution. However, we limit our analysis to Java systems due to the language support of the selected smell-detection tool. While this procedure reduces noise introduced by low-quality or non-representative repositories, it also limits the extent to which our results generalize to non-Java or less popular projects.
A threat to construct validity is the discrepancy between tool-detected code smells and the design problems perceived by developers. While subjective judgments can never be fully captured, we reduce this discrepancy by using a tool that faithfully implements the widely adopted smell definitions of Lanza and Marinescu (Lanza and Marinescu, 2007). Because these definitions underpin many refactoring guidelines, smell-detection tools, and code-smell studies, they approximate a community-level consensus among experienced developers. As a result, the detector’s output is more likely to align with their expectations than tools that rely on opaque or non-standard detection rules. Nevertheless, individual developers may still disagree with specific detections depending on domain knowledge, design intent, or business logic, and our findings should be interpreted with this limitation in mind.
@@
A further threat to construct validity is that we do not include Duplicated Code (i.e., multiple similar or identical code fragments that appear in different locations (Roy and Cordy, 2007)) as a smell in this study. Our theory focuses on ripple effects that propagate along explicit static dependencies between classes, whereas changes related to duplicated code can spread through clone similarity and clone-management practices (e.g., copy-paste) without explicit dependency links. Consequently, clone-related propagation mechanisms may not be captured by our dependency-based measurements, and our findings should be interpreted as applying to dependency-driven rather than clone-driven change propagation.

### v2

@@ 3.2.4. RQ4
We speculate that this situation may further influence stability. This case arises when the focal class and an efferent neighbor both contain code smells (i.e., efferent code smell coupling), and the two smell instances are directly connected by a static dependency where the focal class is the source artifact. In such cases, when a smell in the efferent neighbor changes, the dependency makes it more likely that developers must also change the corresponding smell in the focal class. As a result, developers are more likely to need to modify smell instances in two classes at once. In contrast, in efferent code smell coupling without a direct static dependency between the smell instances, the probability of needing to change smells in both classes at once is lower. Additionally, since interacting code smells are associated with more static dependencies between them (undefm, ), there are more routes for changes to ripple into the focal class, making it more likely to be changed more often and more extensively. Hence, our RQ4 and its hypotheses are:
@@ 3.3. Independent Variables (IVs)
The rationale for measuring the number and variety of code smells is that higher values are often associated with degraded maintainability, particularly reduced stability (undefg, ; undefb, ). The rationale for measuring interaction intensity is that higher intensity reflects a greater number of dependencies connecting smelly artifacts, which can harm comprehensibility and compound a developer’s ability to effectively perform modifications (undefl, ).
@@ 3.5. Control Variables (CVs)
The rationale for including these controls is to reduce confounding. Larger classes inherently have a higher probability of being changed (undefd, ), and classes with more efferent neighbors are also more likely to change because modifications to any of their efferent neighbors may trigger changes in the focal class (undefw, ; undefx, ).
@@ 3.7. System Selection
We will conduct our study on top-starred open-source projects from GitHub. The rationale for using top-starred projects is to prioritize the most popular and widely adopted projects (undefb, ). By ranking projects in descending order of the number of stars, we will select the first 100 projects that meet the following inclusion criteria (IC), following recommendations from Kalliamvakou et al. (undefan, ): IC1. More than 100 forks (indicates meaningful developer interest). IC2. At least 20 contributors (to avoid personal or low-activity projects). IC3. At least one year of commit history with at least 50 commits within that one-year period (to ensure meaningful development activity, approximately one commit per week). IC4. More than 80% of the code written in Java (due to the smell detection tool limitation). IC5. Not an educational project (educational projects may not represent real software development processes).
@@ 3.8. Data Analysis
It is reported that most software engineering metrics do not follow a normal distribution (undefao, ). Therefore, we assume our IVs, DVs, and CVs are non-normal. We use statistical methods that do not assume normality and apply log transformations when appropriate. Due to the large-scale nature of our dataset, normality tests become overly sensitive; therefore, even though we assume non-normality, we verify practical deviations by examining quantile-quantile plots.
@@
We analyze our DVs using generalized linear models (GLMs). GLMs extend ordinary linear regression to non-normal outcomes by using an appropriate distribution and a link function to relate predictors to the expected outcome (undefap, ). In empirical software engineering, regression-based models such as GLMs are widely used to model a variety of software engineering data (undefaq, ; undefar, ). In our analysis, for each hypothesis, we fit a separate GLM with the corresponding DV (either ChF or ChS) as the response and the IV and CVs specified for that hypothesis in Table 2 as predictors.
@@
Both DVs are count data (i.e., non-negative integers arising from counting) and are modeled using negative binomial GLMs with a log link. Poisson and negative binomial GLMs are both standard for count data, but Poisson assumes equidispersion (the mean equals the variance) (undefap, ). Because previous work suggests software engineering count measures can be over-dispersed relative to this assumption, we adopt the negative binomial model for its flexibility (undefas, ). For completeness, we also fit Poisson GLMs and compute the dispersion statistic to confirm the presence of overdispersion.
@@
Since we assume non-normality, we apply log transformations to our count-based IVs (ClSize, #EffNei, #SmellFoc, #SmellEff, VarSmellFoc, VarSmellEff) before modeling, to reduce the influence of extreme values and obtain smoother relationships (undefat, ). We do not transform the DVs, as they are modeled directly using appropriate GLMs.
@@
A random effect is a model component that captures group-level variability, and it is increasingly used in empirical software engineering (undefau, ; undefav, ; undefaw, ). We include a project-level random intercept in all models to account for systematic differences between projects, such as variation in age, development activity, commit practices, and team size. This adjustment ensures that comparisons between classes are not biased by project-specific factors while allowing the estimated effects of code smells to generalize across projects (undefax, ). Using a random intercept rather than fixed effects is appropriate given the large number of projects and our aim to estimate population-level effects that generalize beyond the specific projects, rather than fitting a separate parameter for each project (undefau, ).
@@
For each fitted model, we will report the regression coefficient ($\beta$) for each IV together with its standard error, 95% confidence interval, and $p$-value (computed using Wald tests for the regression coefficients; one-sided as specified by each directional hypothesis), which collectively describe the estimated direction of the effect, its precision, and its statistical significance. To convey the practical magnitude of effects (i.e., effect size), we will report Incidence Rate Ratios (IRRs) (undefay, ) and Average Marginal Effects (AMEs) (undefaz, ). These translate the estimated coefficients into interpretable changes in complementary ways: IRRs present effects as percentage (multiplicative) changes, while AMEs present effects as average raw-unit (additive) changes. Model performance will be summarized using the log-likelihood (undefap, ) and McFadden’s pseudo-$R^{2}$ (undefaaa, ), both of which provide standard goodness-of-fit evaluations for GLMs. Together, these reporting elements ensure a comprehensive assessment of model adequacy.
@@ 3.9. Implementation
Projects will be filtered (according to the first four inclusion criteria discussed in Section 3.7), selected, and downloaded automatically using the GitHub API (undefaab, ). Inclusion Criterion 5 (i.e., not an educational project) will be enforced manually by inspecting each project repository, primarily the project documentation (e.g., the README). Code smells will be detected using JSpIRIT (undefaac, ; undefaad, ). The rationale for choosing this tool is that it can detect the ten code smells we will study, and it strictly follows the definitions and detection strategies provided by Lanza and Marinescu (undefn, ), ensuring transparency. Static dependencies, efferent neighbors of the focal class, and class size will be detected and measured using the static analysis tool CodeQL (undefaae, ). This information will be collectively analysed using Python to obtain all IVs and CVs. Git commands (undefaaf, ) will be used to analyse commit history and extract information such as diffs, and RefactoringMiner (undefaag, ) will be used to identify class refactorings, including class rename, move, split, and merge. Python will automate the process and compute the DVs for each included class. Finally, R will be used to fit the GLMs via the glmmTMB package (undefaah, ) and to obtain all statistical outputs.
@@ 4. Threats to Validity
A further threat to internal validity concerns our project selection process. Following the recommendations of Kalliamvakou et al. (undefan, ), we exclude personal, inactive, and educational repositories and focus on widely adopted projects to maximize insights into real code evolution. However, we limit our analysis to Java systems due to the language support of the selected smell-detection tool. While this procedure reduces noise introduced by low-quality or non-representative repositories, it also limits the extent to which our results generalize to non-Java or less popular projects.
A threat to construct validity is the discrepancy between tool-detected code smells and the design problems perceived by developers. While subjective judgments can never be fully captured, we reduce this discrepancy by using a tool that faithfully implements the widely adopted smell definitions of Lanza and Marinescu (undefn, ). Because these definitions underpin many refactoring guidelines, smell-detection tools, and code-smell studies, they approximate a community-level consensus among experienced developers. As a result, the detector’s output is more likely to align with their expectations than tools that rely on opaque or non-standard detection rules. Nevertheless, individual developers may still disagree with specific detections depending on domain knowledge, design intent, or business logic, and our findings should be interpreted with this limitation in mind.
@@
A further threat to construct validity is that we do not include Duplicated Code (i.e., multiple similar or identical code fragments that appear in different locations (undefaai, )) as a smell in this study. Our theory focuses on ripple effects that propagate along explicit static dependencies between classes, whereas changes related to duplicated code can spread through clone similarity and clone-management practices (e.g., copy-paste) without explicit dependency links. Consequently, clone-related propagation mechanisms may not be captured by our dependency-based measurements, and our findings should be interpreted as applying to dependency-driven rather than clone-driven change propagation.